R3al:

Factor: 5x2 – 15x – 20. (a) 5(x-4)(x+1), (b) -2(x-4)(x+5), (c) -5(x+4)(x-1), (d) 5(x+4)(x+1).

MasterEmoji09:

what do u think it is

R3al:

Oh, Sorry wrong one.

jhonyy9:

factoriz out the 5 and get 5(x^2 -3x -4) so in this case just you need getting 2 numbers with sum equal -3 and with product equal -4 a+b=-3 a*b= -4 when you get these so this above wrote quadratic will be in this form 5(x^2 +Sx +P) where S = a +b P = a*b hope helped understandably easy